MILLIMETER-WAVE SPECTROSCOPY OF H_{2}-HCl CLUSTER

\underline{MASAZUMI ISHIGURO}, K. HARADA, K. TANAKA AND T. TANAKA, Devision of Chemistry, Graduate School of Science, Kyushu University, Fukuoka, 812-8581 Japan.

The pure rotational spectrum of the (p-H_{2})- and (o-H_{2})-HCl clusters for both ^{35}Cl and ^{37}Cl species have been recorded by using pulsed supersonic jet technique and millimeter-wave direct absorption spectroscopy in the 67-166 GHz region. These spectra were searched by using prediction based on the result of the infrared spectrum of the Cl-H stretching band.

a

For (p-H_{2})-HCl, two rotational transitions, J = 3 − 2, 4 − 3, in the Σ_{0} ground state were obseved. For (o-H_{2})-HCl, four rotational transitions split into l-type doubling, J = 2 − 1 ∼ 5 − 4, in the Π_{0} ground state were observed. All of spectra were split into hyperfine structure due to Cl nucleus. The intensity of (p-H_{2})-HCl were approximately 1 / 20 that of (o-H_{2})-HCl. The rotatonal constant, B_{0}, for the (p-H_{2})-H^{35}Cl and (o-H_{2})-H^{35}Cl were determined to be 16627.2089(20) and 16841.5120(54) MHz, respectively. The analysis of hyperfine structure and the search for rovibrational transions of the vdW bending mode are now in progress.

aD. T. Anderson et al., Chem. Phys. \underline{239}, 253-269 (1998)